40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 1645     Entry time: Wed Jun 3 03:22:16 2009
Author: pete 
Type: Update 
Category: Locking 
Subject: DD handoff 

Rana, Alberto, Pete

We have the DD handoff nominally working.  Sometimes, increasing the SRC gain at the end makes MICH get unstable.  This could be due to a non-diagonal term in the matrix, or possibly because the DRM locks in a funky mode sometimes. 

To get the DD handoff working, first we tuned demod phases in order to zero the offsets in the PD signals handed-off-to.  Based on transer function measurements, I set the PRC PD6_I element to 0.1, and set the PD8_I signal to 0, since it didn't seem to be contributing much.  We also commented out the MICH gain increase at the end of the DD_handoff script.

It could still be more stable, but it seems to work most of the time.

 

 

ELOG V3.1.3-